OpenStudy (anonymous):

Solve: -4(x + 10) - 6 = -3(x - 2)

OpenStudy (anonymous):

-4x-40-6=-3x+6 -4x-46=-3x+6 -x=52 x=-52

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

if i dont get my work accomplished than i might fail, and that would suck

OpenStudy (anonymous):

np(: simply first then regroup the x terms to one side and constants to other side, so that it would be easy to see what you're solving for(:

OpenStudy (anonymous):

alright cool